OpenStudy (anonymous):

Input the equation of the given line in standard form. The line with m = and the point (-2, -3).

OpenStudy (anonymous):

m=-1/2

OpenStudy (anonymous):

standard from looks like: Ax + By = C that is where we need to end up, but we will start at: y = mx+b can you fill in what y, m, x should be?

OpenStudy (anonymous):

with that I get b=-4? how do i solve for Ax+By=C??

OpenStudy (anonymous):

correct so first plug back in m and b into y=mx+b then, isolate the 'b' term so its all by its self